# Quantum One-Time Programs, Revisited

### v1

@@ 6.1.3 Subspace Coset States and Computational Direct-Product Hardness
Definition 6.5 (Subspace Coset States, [CLLZ21b, VZ21]).
@@
We then present the computational version of the direct product hardness property of Theorem 3.9, by combining two theorems from [CLLZ21b] and [CHV23].
@@
Theorem 6.7 (Computational Direct Product Hardness, [CLLZ21b, CHV23]).
@@ 8.2 One-Time NIZK Proofs
The common reference string $\mathsf{CRS}$ consists of the two obfuscated programs.
It also outputs {|$\{\ket{A_{i}}\}_{i\in[n]}$ as the one-time proof token.
@@ 8.4 One-time programs for Verifiable Functions Imply Quantum Money
The correctness of this quantum money scheme follows from the correctness of the one-time program $\mathsf{OTP}(f)$ and the verification procedure $\mathsf{Ver}_{\mathcal{F}}$.
We now prove security. Suppose for contradiction there exists an adversary $A$ that wins the quantum money security game with non-negligible probability, so that it outputs a joint state $\rho_{1,2}$ on two registers $\rho_{1}$ and $\rho_{2}$ such that both verification checks pass. Then, there exists an adversary $B$ that breaks the one-time program security of $\mathsf{OTP}(f)$: given $\mathsf{OTP}(f)$ and $vk_{f}$ as auxiliary information, $B$ runs $A$ on the same input, $A(\mathsf{OTP}(f),vk_{f})$. Suppose that $A$ outputs (a joint state on) two registers $\rho_{1},\rho_{2}$. Then, running the verification procedure on both these registers results in two samples $(x_{1},y_{1})$ and $(x_{2},y_{2})$, breaking the one-time program security of $\mathsf{OTP}$. ∎
@@ A.1 Pairwise Independent and Highly Random Functions
    Note that $\mathbb{I}-\sum_{y^{\prime}\in\mathcal{Y}}\ket{F_{x,r^{\prime},y^{\prime}}}\bra{F_{x,r^{\prime},y^{\prime}}}$ | is a projector because the states (|$\left(\ket{F_{x,r^{\prime},y^{\prime}}}\right)_{y\in\mathcal{Y}}$ are mutually orthogonal. Therefore,
@@
    $M_{x}^{\dagger}\cdot M_{x}$ has a simple spectral decomposition.
    First, let |$\ket{\Phi_{0}}$ be a column vector with entries $\left(\sqrt{\frac{\Pr_{f}[f(x,r)=y]}{|\mathcal{R}|}}\right)_{(r,y)\in\mathcal{R}\times\mathcal{Y}}$. Note that
@@
    Then for any $(r,y),(r^{\prime},y^{\prime})$, the $(r,y),(r^{\prime},y^{\prime})$-th entry of |$\ket{\Phi_{0}}\bra{\Phi_{0}}$ | is $\frac{1}{|\mathcal{R}|}\cdot\sqrt{\Pr_{f}[f(x,r)=y]}\cdot\sqrt{\Pr_{f}[f(x,r^{\prime})=y^{\prime}]}$.
    Second, let $N_{x}$ be the following matrix. If $r\neq r^{\prime}$, then $(r,y),(r^{\prime}y^{\prime})$-th entry is 00. If $r=r^{\prime}$, then the $(r,y),(r^{\prime}y^{\prime})$-th entry is $\frac{1}{|\mathcal{R}|}\cdot\sqrt{\Pr_{f}[f(x,r)=y]}\cdot\sqrt{\Pr_{f}[f(x,r^{\prime})=y^{\prime}]}$. Note that $N_{x}$ is positive-semi-definite.
    Third,
@@
    $\|(\mathbb{I}-\ket{F_{\emptyset}}\bra{F_{\emptyset}})\cdot M_{x}\cdot\ket{\Phi}\|_{2}$ $\leq|\beta|\cdot\|(\mathbb{I}-\ket{F_{\emptyset}}\bra{F_{\emptyset}})\|_{2}\cdot\|M_{x}\cdot\ket{\Phi_{1}}\|$
@@ B.2 Security Proof
Appendix C Recording Queries for High Entropy Functions in the Single-Effective-Query Model
The following claim shows that if $f$ is highly random, then SEQ queries to $f$ can be detected from the oracle register with high probability. The state of $\mathcal{H}_{x}$ after querying $f(x,h(x))$ on $x$ will be $\rho_{x}$, which has negligible fidelity with |$\ket{R}\bra{R}$ |, the state before the query.
Claim C.1.
For a given $x\in\mathcal{X}$: if for all $y\in\mathcal{Y}$, $\Pr_{R\overset{\$}{\leftarrow}\mathcal{R}}[y=f(x,R)]=\mathsf{negl}(\lambda)$, then
$|\bra{R}\rho_{x}\ket{R}|=\mathsf{negl}(\lambda)$
Proof.
$\bra{R}\rho_{x}\ket{R}$ $=\left(\frac{1}{\sqrt{|\mathcal{R}|}}\cdot\sum_{r^{\prime\prime}\in\mathcal{R}}\bra{r^{\prime\prime}}\right)\cdot\left(\frac{1}{|\mathcal{R}|}\cdot\sum_{r,r^{\prime}\in\mathcal{R}}\ket{r}\bra{r^{\prime}}\cdot\mathbbm{1}_{f(x,r)=f(x,r^{\prime})}\right)\cdot\left(\frac{1}{\sqrt{|\mathcal{R}|}}\cdot\sum_{r^{\prime\prime\prime}\in\mathcal{R}}\ket{r^{\prime\prime\prime}}\right)$
$=\frac{1}{|\mathcal{R}|^{2}}\cdot\sum_{r,r^{\prime},r^{\prime\prime},r^{\prime\prime\prime}\in\mathcal{R}}\braket{r^{\prime\prime}}{r}\braket{r^{\prime}}{r^{\prime\prime\prime}}\cdot\mathbbm{1}_{f(x,r)=f(x,r^{\prime})}$
$=\frac{1}{|\mathcal{R}|^{2}}\cdot\sum_{r,r^{\prime}\in\mathcal{R}}\mathbbm{1}_{f(x,r)=f(x,r^{\prime})}$
$=\frac{1}{|\mathcal{R}|^{2}}\cdot\sum_{y\in\mathcal{Y}}|\{r\in\mathcal{R}:y=f(x,r)\}|^{2}$
$=\sum_{y\in\mathcal{Y}}\Pr_{R}[y=f(x,R)]^{2}$
Let $\nu(\lambda)$ be the maximum value of $\Pr_{R}[y=f(x,R)]$, over all $y$-values. Note that $\nu(\lambda)=\mathsf{negl}(\lambda)$. Furthermore,
11 11 $=\sum_{y\in\mathcal{Y}}\Pr_{R}[y=f(x,R)]$
$\nu(\lambda)$ $=\sum_{y\in\mathcal{Y}}\Pr_{R}[y=f(x,R)]\cdot\nu(\lambda)$
$\geq\sum_{y\in\mathcal{Y}}\Pr_{R}[y=f(x,R)]\cdot\Pr_{R}[y=f(x,R)]=\sum_{y\in\mathcal{Y}}\Pr_{R}[y=f(x,R)]^{2}$
$\mathsf{negl}(\lambda)$ $\geq\bra{R}\rho_{x}\ket{R}$
Appendix D More on the Compressed Oracle
D.1 Alternative Representations of the Decompression Function
@@
D.2 Compressed Oracle Chaining
@@
Following almost the same analysis, we will also show that if an adversary has access to a function $F(x)=H(x,G(x))$ and $H$ records $H(x\|y)=z$, then with overwhelming probability $G$ contains a matching entry $G(x)=y$. See Lemma D.4 at the end of this subsection for more details.
Lemma D.1 (Compressed Oracle Chaining).
@@
For basis vectors in the support of $E_{Y1,Z}$, we can write $D_{G}=D_{G}^{\prime}\cup(x,y)$ and $D_{H}=D_{H}^{\prime}\cup(y,z)$, where $D_{G}^{\prime}(x)=\bot$ and $D_{H}^{\prime}(y)=\bot$. $D_{G}^{\prime}$ and $D_{H}^{\prime}$ represent $D_{G}$ and $D_{H}$ with $x$ and $y$ removed, respectively. Furthermore, there does not exist an $x^{\prime}\in\mathcal{X}$ such that $D_{G}^{\prime}(x^{\prime})=y$, since $x$ is the unique preimage of $y$ under $G$. Recall from Section D.1 that the effect of $\mathsf{Decomp}_{G}$ on |$\ket{x,0,D\cup(x,y)}$ is
@@
Putting together D.2 and D.3, the norm after any single query to $H\circ G$ is bounded as
@@
The proof requires changing just a few lines of the proof of Lemma D.1. The first modification is that $U_{H}$ acts on all three registers of $\mathcal{Q}=(\mathcal{Q}_{\mathcal{X}},\mathcal{Q}_{\mathcal{Y}},\mathcal{Q}_{\mathcal{Z}})$, although it only modifies $\mathcal{Q}_{\mathcal{Z}}$. This change does not affect the proof of D.2.
Second, in the proof of D.3, the projectors are slightly modified. $E_{Y,\bot}$ now requires $D_{H}(x\|y)=\bot$, instead of $D_{H}(y)=\bot$. Additionally, $E_{Y1,Z}$ and $E_{Y+,Z}$ are replaced by a single new projector $E_{Y,Z}$, which projects onto states |$\ket{a}\otimes\ket{x,0,u}\otimes\ket{D_{G},D_{H}}$ where (1) $D_{G}$ and $D_{H}$ satisfy $E$, (2) $D_{G}(x)=y$ for some $y$, and (3) $D_{H}(x\|y)=z$ for some $z$. The fourth condition of uniqueness is no longer necessary because $x\|y$ fully specifies $x$. The analysis of $\left\|\overline{E}\cdot\mathsf{Decomp}_{G}\cdot E_{Y,Z}\ket{\psi^{\prime}}\right\|$ is almost exactly the same as that of $\left\|\overline{E}\cdot\mathsf{Decomp}_{G}\cdot E_{Y1,Z}\ket{\psi^{\prime}}\right\|$ in the original proof, with the minor syntactic change that $D_{H}$ contains a tuple $(x\|y,z)$ instead of $(y,z)$. ∎
D.3 Knowledge of Preimage for Expanding Random Oracles
Here we show that if a random oracle $G$ is sufficiently expanding and an adversary making polynomially many queries to $G$ knows elements $y$ that appear in the image of $G$, then $G$’s compressed database should also contain a corresponding entry.
Lemma D.5.
Let $G:\mathcal{X}\rightarrow\mathcal{Y}$ be a random function where $|\mathcal{X}|<|\mathcal{Y}|$. Consider an oracle algorithm $A$ makes $q$ of queries to $G$, then outputs a set of $k$ values $\mathbf{y}=(y_{1},\dots,y_{k})$. Let $p$ be the probability that for every $i$, there exists an $x_{i}\in\mathcal{X}$ such that $G(x_{i})=y_{i}$.
Now consider running the same experiment where $G$ is instead implemented as a compressed oracle, and measuring its database register after $A$ outputs to obtain $D$. Let $p^{\prime}$ be the probability that for every $i$, there exists an $x_{i}\in\mathcal{X}$ such that $D(x_{i})=y_{i}$. If $k$ and $q$ are $\mathsf{poly}(\lambda)$ and $|\mathcal{X}|^{k}/|\mathcal{Y}|=\mathsf{negl}(\lambda)$, then [^17] We remark that the reliance on the number of queries is unlikely to be tight. A tighter bound might be achieved by performing a direct computation of the effects of querying $G$ on every $x\in\mathcal{X}$ at the end of the experiment.

### v2

@@ 6.1.3 Subspace Coset States and Computational Direct-Product Hardness
Definition 6.5 (Subspace Coset States, [CLLZ21a, VZ21]).
@@
We then present the computational version of the direct product hardness property of Theorem 3.9, by combining two theorems from [CLLZ21a] and [CHV23].
@@
Theorem 6.7 (Computational Direct Product Hardness, [CLLZ21a, CHV23]).
@@ 8.2 One-Time NIZK Proofs
We defer the regular soundness and zero-knowledge definition of NIZK tp Appendix D.
@@
The common reference string $\mathsf{CRS}$ consists of the two obfuscated programs and a master secret key $\mathsf{msk}=\{A_{i}\}_{i\in[n]}$.
$\mathsf{Delegate}$: takes in a master secret key $\mathsf{msk}=\{A_{i}\}_{i\in[n]}$ and outputs {|$\{\ket{A_{i}}\}_{i\in[n]}$ as the one-time proof token.
@@
We then prove the following statement:
Theorem 8.3.
Assuming LWE and subexponentially secure iO, there exists one-time NIZK proofs satisfying the above definition.
@@ 8.4 One-time programs for Verifiable Functions Imply Quantum Money
In this section, we show that our definition of one-time (sampling) program for publicly verifiable functions implies public -key quantum money.
@@
The correctness of this quantum money scheme follows from the correctness of the one-time program $\mathsf{OTP}(f)$, the verification procedure $\mathsf{Ver}_{\mathcal{F}}$ as well as the gentle measurement lemma Lemma 3.2.
We now prove security. Suppose for contradiction there exists an adversary $A$ that wins the quantum money security game with non-negligible probability, so that it outputs a joint state $\rho_{1,2}$ on two registers $\rho_{1}$ and $\rho_{2}$ such that both verification checks pass. Then, there exists an adversary $B$ that breaks the one-time program security of $\mathsf{OTP}(f)$: given $\mathsf{OTP}(f)$ and $vk_{f}$ as auxiliary information, $B$ runs $A$ on the same input, $A(\mathsf{OTP}(f),vk_{f})$. Suppose that $A$ outputs (a joint state on) two registers $\rho_{1},\rho_{2}$. Then, since the verification procedure verifies by sampling a random input to test the functionality on each state, running the verification procedure on both these registers results in two samples $(x_{1},y_{1})$ and $(x_{2},y_{2})$ with non-negligible probability, breaking the one-time program security of $\mathsf{OTP}$. ∎
@@ A.1 Pairwise Independent and Highly Random Functions
    Note that $\mathbb{I}-\sum_{y^{\prime}\in\mathcal{Y}}\ket{F_{x,r^{\prime},y^{\prime}}}\bra{F_{x,r^{\prime},y^{\prime}}}$ | is a projector because the states (|$\left(\ket{F_{x,r^{\prime},y^{\prime}}}\right)_{y^{\prime}\in\mathcal{Y}}$ are mutually orthogonal. Therefore,
@@
    =$=\begin{cases}\frac{1}{|\mathcal{R}|},&(r,y)=(r^{\prime},y^{\prime})\\ 0,&r=r^{\prime}\land y\neq y^{\prime}\\ \frac{1}{|\mathcal{R}|}\cdot\sqrt{\Pr_{f}[f(x,r)=y]}\cdot\sqrt{\Pr_{f}[f(x,r^{\prime})=y^{\prime}]},&r\neq r^{\prime}\end{cases}$
    by lemma A.7.
@@
    Let us write $M_{x}^{\dagger}\cdot M_{x}$ as a linear combination of PSD matrices.
    First, let |$\ket{\Phi_{0}}$ be a column vector such that the $(r,y)$-th entry is $\sqrt{\frac{\Pr_{f}[f(x,r)=y]}{|\mathcal{R}|}}$. Additionally, let |$\ket{\Phi_{0,r}}$ be a column vector such that the $(r^{\prime},y)$-th entry is 00 if $r\neq r^{\prime}$, and $\sqrt{\Pr_{f}[f(x,r)=y]}$ if $r=r^{\prime}$.
    Note that these vectors have unit norm:
@@
    |$\|\ket{\Phi_{0,r}}\|_{2}^{2}$ $=\sum_{y\in\mathcal{Y}}\Pr_{f}[f(x,r)=y]=1$
    Second,
    let $\text{let }N_{x}$ $=\sum_{r\in\mathcal{R}}\frac{1}{|\mathcal{R}|}\cdot\ket{\Phi_{0,r}}\bra{\Phi_{0,r}}$ |
    Note that (|$\left(\ket{\Phi_{0,r}}\right)_{r\in\mathcal{R}}$ are mutually orthogonal, so $N_{x}$ is positive semi-definite.
    Third, we claim that $M_{x}^{\dagger}\cdot M_{x}$ can be written in the following form:
@@
    The $(r,y),(r^{\prime},y^{\prime})$-th entry of $\left(\frac{1}{|\mathcal{R}|}\cdot\mathbb{I}+\ket{\Phi_{0}}\bra{\Phi_{0}}-N_{x}\right)$ is given below. As shorthand, we let $p_{x,r,y}=\Pr_{f}[f(x,r)=y]$.
    $\left(\frac{1}{|\mathcal{R}|}\cdot\mathbb{I}+\ket{\Phi_{0}}\bra{\Phi_{0}}-N_{x}\right)_{(r,y),(r^{\prime},y^{\prime})}$
    =$\quad\quad=\begin{cases}\frac{1}{|\mathcal{R}|}\left(1+p_{x,r,y}-p_{x,r,y}\right),&(r,y)=(r^{\prime},y^{\prime})\\ \frac{1}{|\mathcal{R}|}\left(0+\sqrt{p_{x,r,y}}\cdot\sqrt{p_{x,r^{\prime},y^{\prime}}}-\sqrt{p_{x,r,y}}\cdot\sqrt{p_{x,r^{\prime},y^{\prime}}}\right),&r=r^{\prime}\land y\neq y^{\prime}\\ \frac{1}{|\mathcal{R}|}\left(0+\sqrt{p_{x,r,y}}\cdot\sqrt{p_{x,r^{\prime},y^{\prime}}}-0\right),&r\neq r^{\prime}\end{cases}$
    =$\quad\quad=\begin{cases}\frac{1}{|\mathcal{R}|},&(r,y)=(r^{\prime},y^{\prime})\\ 0,&r=r^{\prime}\land y\neq y^{\prime}\\ \frac{1}{|\mathcal{R}|}\cdot\sqrt{p_{x,r,y}}\cdot\sqrt{p_{x,r^{\prime},y^{\prime}}},&r\neq r^{\prime}\end{cases}$
    $\quad\quad=(M_{x}^{\dagger}\cdot M_{x})_{(r,y),(r^{\prime},y^{\prime})}$
    Therefore, $M_{x}^{\dagger}\cdot M_{x}=\frac{1}{|\mathcal{R}|}\cdot\mathbb{I}+\ket{\Phi_{0}}\bra{\Phi_{0}}-N_{x}$
@@
    We used the fact that $\bra{\Phi_{1}}\cdot N_{x}\cdot\ket{\Phi_{1}}\geq 0$ because $N_{x}$ is PSD.
@@
    $\|(\mathbb{I}-\ket{F_{\emptyset}}\bra{F_{\emptyset}})\cdot M_{x}\cdot\ket{\Phi}\|_{2}$ $\leq|\beta|\cdot\|(\mathbb{I}-\ket{F_{\emptyset}}\bra{F_{\emptyset}})\|_{2}\cdot\|M_{x}\cdot\ket{\Phi_{1}}\|_{2}$
@@ B.2 Security Proof
Appendix C More on the Compressed Oracle
C.1 Alternative Representations of the Decompression Function
@@
C.2 Compressed Oracle Chaining
@@
Following almost the same analysis, we will also show that if an adversary has access to a function $F(x)=H(x,G(x))$ and $H$ records $H(x\|y)=z$, then with overwhelming probability $G$ contains a matching entry $G(x)=y$. See Lemma C.4 at the end of this subsection for more details.
Lemma C.1 (Compressed Oracle Chaining).
@@
For basis vectors in the support of $E_{Y1,Z}$, we can write $D_{G}=D_{G}^{\prime}\cup(x,y)$ and $D_{H}=D_{H}^{\prime}\cup(y,z)$, where $D_{G}^{\prime}(x)=\bot$ and $D_{H}^{\prime}(y)=\bot$. $D_{G}^{\prime}$ and $D_{H}^{\prime}$ represent $D_{G}$ and $D_{H}$ with $x$ and $y$ removed, respectively. Furthermore, there does not exist an $x^{\prime}\in\mathcal{X}$ such that $D_{G}^{\prime}(x^{\prime})=y$, since $x$ is the unique preimage of $y$ under $G$. Recall from Section C.1 that the effect of $\mathsf{Decomp}_{G}$ on |$\ket{x,0,D\cup(x,y)}$ is
@@
Putting together C.2 and C.3, the norm after any single query to $H\circ G$ is bounded as
@@
The proof requires changing just a few lines of the proof of Lemma C.1. The first modification is that $U_{H}$ acts on all three registers of $\mathcal{Q}=(\mathcal{Q}_{\mathcal{X}},\mathcal{Q}_{\mathcal{Y}},\mathcal{Q}_{\mathcal{Z}})$, although it only modifies $\mathcal{Q}_{\mathcal{Z}}$. This change does not affect the proof of C.2.
Second, in the proof of C.3, the projectors are slightly modified. $E_{Y,\bot}$ now requires $D_{H}(x\|y)=\bot$, instead of $D_{H}(y)=\bot$. Additionally, $E_{Y1,Z}$ and $E_{Y+,Z}$ are replaced by a single new projector $E_{Y,Z}$, which projects onto states |$\ket{a}\otimes\ket{x,0,u}\otimes\ket{D_{G},D_{H}}$ where (1) $D_{G}$ and $D_{H}$ satisfy $E$, (2) $D_{G}(x)=y$ for some $y$, and (3) $D_{H}(x\|y)=z$ for some $z$. The fourth condition of uniqueness is no longer necessary because $x\|y$ fully specifies $x$. The analysis of $\left\|\overline{E}\cdot\mathsf{Decomp}_{G}\cdot E_{Y,Z}\ket{\psi^{\prime}}\right\|$ is almost exactly the same as that of $\left\|\overline{E}\cdot\mathsf{Decomp}_{G}\cdot E_{Y1,Z}\ket{\psi^{\prime}}\right\|$ in the original proof, with the minor syntactic change that $D_{H}$ contains a tuple $(x\|y,z)$ instead of $(y,z)$. ∎
C.3 Knowledge of Preimage for Expanding Random Oracles
Here we show that if a random oracle $G$ is sufficiently expanding and an adversary making polynomially many queries to $G$ knows elements $y$ that appear in the image of $G$, then $G$’s compressed database should also contain a corresponding entry. In fact, we will show a more general statement which takes into account an adversary which knows partial preimages as well. In the case where $\mathcal{X}_{1}=\emptyset$ in the following lemma, this corresponds to an adversary simply finding elements from the range.
Lemma C.5.
Let $G:\mathcal{X}_{1}\times\mathcal{X}_{2}\rightarrow\mathcal{Y}$ be a random function where $|\mathcal{X}_{2}|<|\mathcal{Y}|$. Consider an oracle algorithm $A$ makes $q$ of queries to $G$, then outputs two vector of $k$ values $\mathbf{x^{(1)}}=\left(x_{1}^{(1)}\dots,x_{k}^{(1)}\right)$ and $\mathbf{y}=(y_{1},\dots,y_{k})$. Let $p$ be the probability that for every $i$, there exists an $x_{i}^{(2)}\in\mathcal{X}$ such that $G\left(x_{i}^{(1)},x_{i}^{(2)}\right)=y_{i}$.
Now consider running the same experiment where $G$ is instead implemented as a compressed oracle, and measuring its database register after $A$ outputs to obtain $D$. Let $p^{\prime}$ be the probability that for every $i$, there exists an $x_{i}^{(2)}\in\mathcal{X}_{2}$ such that $D\left(x_{i}^{(1)}\|x_{i}^{(2)}\right)=y_{i}$. If $k$ and $q$ are $\mathsf{poly}(\lambda)$ and $|\mathcal{X}_{2}|^{k}/|\mathcal{Y}|=\mathsf{negl}(\lambda)$, then [^17] We remark that the reliance on the number of queries is unlikely to be tight. A tighter bound might be achieved by performing a direct computation of the effects of querying $G$ on every $x\in\mathcal{X}$ at the end of the experiment.